OpenStudy (lgbasallote):

i think im beginning to get the hang of trig substitution \m/ thank you @apoorvk and @Mimi_x3 :DDD but another wild question...when and how to use hyperbolic functions in trig sub?

OpenStudy (lgbasallote):

should i write a problem?

OpenStudy (lgbasallote):

Use the hyperbolic substitution x = a sinh t to show that \(\huge \int \frac{dx}{\sqrt{x^2 + a^2}} = \sinh^{-1} (\frac{x}{a}) + C\)

OpenStudy (apoorvk):

Sorry mate, I have no idea about hyperbolic substitutions. :( Though the answer is the same using simple substitution (i.e. sin^(-1)(x/a) +c )

OpenStudy (apoorvk):

use x=tanA

OpenStudy (lgbasallote):

oh okay :DDD anyway...back to learning trig sub